Chapter 8

Limit Theorems. Laws of Large Numbers. Central Limit Theorems

[19/100%]
Show
LaTeX
Problem 251

If the independent random variables X1,…,Xn,…X_{1}, \ldots , X_{n}, \ldots satisfy the condition

Var⁡[Xi]≤c<∞,i=1,2,…, \operatorname {Var}\left[X_{i}\right] \leq c < \infty , \qquad i = 1, 2, \ldots ,

then the SLLN holds.

?
Problem 252

In a sequence of random variables X1,…,Xn,…X_{1}, \ldots , X_{n}, \ldots suppose that XkX_{k} depends only on Xk−1,Xk+1X_{k-1}, X_{k+1}, but that it is independent of all the other random variables (k=2,3,…k = 2, 3, \ldots). Show that if Var⁡[Xi]≤N<∞\operatorname {Var}\left[X_{i}\right] \leq N < \infty (i=1,2,…i = 1, 2, \ldots), then the WLLN holds.

?
Problem 253

If for every nn, Var⁡[Xi]≤c<∞\operatorname {Var}\left[X_{i}\right] \leq c < \infty and Cov⁡[Xi,Xj]<0\operatorname {Cov}\left[ X_{i}, X_{j} \right] < 0 (i,j=1,2,…,ni, j = 1, 2, \ldots , n), then the WLLN holds.

?
Problem 254

Let {Xn}\left\{ X_{n}\right\} be a sequence of random variables so that Var⁡[Xi]≤c<∞\operatorname {Var}\left[X_{i}\right] \leq c < \infty (i=1,2,…i = 1, 2, \ldots) and Cov⁡[Xi,Xj]→0\operatorname {Cov}\left[ X_{i}, X_{j} \right] \to 0 when ∣i−j∣→∞\left|i - j\right| \to \infty then the WLLN holds (Theorem of Bernstein).

?
Problem 255

If Xn→LXX_{n} \xrightarrow {L} X, Yn→PcY_{n} \xrightarrow {P} c, then

?
(a)

Xn+Yn→LX+cX_{n} + Y_{n} \xrightarrow {L} X + c;

(b)

XnYn→LcXX_{n} Y_{n} \xrightarrow {L} cX;

(c)

Xn/Yn→LX/cX_{n} / Y_{n} \xrightarrow {L} X / c.

Problem 256

Let ξn=(X1n,X2n,…,Xkn)′\xi_{n} = (X_{1n}, X_{2n}, \ldots , X_{kn})' (n=1,2,…n = 1, 2, \ldots) be a sequence of random vectors and ξ=(X1,X2,…,Xk)′\xi = (X_{1}, X_{2}, \ldots , X_{k})' a random vector with distribution function F(x1,…,xk)F(x_{1}, \ldots , x_{k}). If

∑i=1kciXin→L∑i=1kciXi \sum _{i=1}^{k} c_{i} X_{in} \xrightarrow {L} \sum _{i=1}^{k} c_{i} X_{i}

for all constants c1,…,ckc_{1}, \ldots , c_{k}, then the joint distribution function FξnF_{\xi_{n}} of X1n,…,XknX_{1n}, \ldots , X_{kn} has a limit and

Fξn(x1,…,xk)→F(x1,…,xk). F_{\xi _{n}}(x_{1}, \ldots , x_{k}) \to F(x_{1}, \ldots , x_{k}).
?
Problem 257

Let {Xn}\left\{ X_{n}\right\} be a sequence of independent random variables uniformly bounded, that is, there is a constant MM such that P(Xk≤M)=1\mathbb {P}\left(X_{k} \leq M\right) = 1 for all k=1,2,…k = 1, 2, \ldots, and suppose that Var⁡[Xk]≠0\operatorname {Var}\left[X_{k}\right] \neq 0 for every kk. If sn2→∞s_{n}^{2} \to \infty, then the CLT holds (cf. (8.13)).

Formula (8.13), Lyapunov's CLT: let X1,…,Xn,…X_{1}, \ldots , X_{n}, \ldots be a sequence of independent random variables with E[Xi]=μi\mathbb {E}\left[X_{i}\right] = \mu_{i}, Var⁡[Xi]=σi2\operatorname {Var}\left[X_{i}\right] = \sigma_{i}^{2}, βi=E[∣Xi−μi∣3]\beta_{i} = \mathbb {E}\left[\left|X_{i} - \mu_{i}\right|^{3}\right], and set sn=(∑i=1nσi2)1/2s_{n} = \left(\sum_{i=1}^{n} \sigma_{i}^{2}\right)^{1/2}, Bn=(∑i=1nβi)1/3B_{n} = \left(\sum_{i=1}^{n} \beta_{i}\right)^{1/3}; if Bn/sn→0B_{n}/s_{n} \to 0 as n→∞n \to \infty, then

Sn∗=∑i=1n(Xi−μi)sn→n→∞LN(0,1). S_{n}^{*} = \frac{\sum _{i=1}^{n} (X_{i} - \mu _{i})}{s_{n}} \xrightarrow [n \to \infty ]{L} N(0, 1).
?
Problem 258

Show that

lim⁡n→∞e−n∑k=0nnkk!=12. \lim _{n \to \infty } e^{-n} \sum _{k=0}^{n} \frac{n^{k}}{k!} = \frac{1}{2}.
?
Problem 259

If XX and YY are independent Poisson variables with parameters λ1\lambda_{1} and λ2\lambda_{2}, respectively, show that

X−Y−(λ1−λ2)(X+Y)1/2→λ1→∞λ2→∞LN(0,1). \frac{X - Y - (\lambda _{1} - \lambda _{2})}{(X+Y)^{1/2}} \xrightarrow [\substack {\lambda _{1} \to \infty \\ \lambda _{2} \to \infty }]{L} N(0, 1).
?
Problem 260

Show that for the sequences {Xn}\left\{ X_{n}\right\} of independent random variables with

?
(a)

P(Xn=±1)=1−2−n2,P(Xn=±2n)=12n+1,n=1,2,…,\displaystyle \mathbb {P}\left(X_{n} = \pm 1\right) = \frac{1 - 2^{-n}}{2}, \qquad \mathbb {P}\left(X_{n} = \pm 2^{n}\right) = \frac{1}{2^{n+1}}, \qquad n = 1, 2, \ldots ,

(b)

P(Xn=±nλ)=1/2\mathbb {P}\left(X_{n} = \pm n^{\lambda }\right) = 1/2,

the CLT holds.

Problem 261

Compare the results given by Chebyshev's inequality and the CLT for the probabilities

P(−k<Sn∗≤k)fork=1,2,3. \mathbb {P}\left(-k < S_{n}^{*} \leq k\right) \qquad \text{for} \qquad k = 1, 2, 3.
?
Problem 262

A variable XX has the Pareto distribution with density

f(x)=24x−4,x≥2. f(x) = 24 x^{-4}, \qquad x \geq 2.

Give the graph of the function g(δ)=P(∣X−μ∣>δ)g(\delta ) = \mathbb {P}\left(\left|X - \mu \right| > \delta \right) where μ=E[X]\mu = \mathbb {E}\left[X\right], and compare this with the upper bound given by Chebyshev's inequality.

?
Problem 263

Given that Var⁡[X]=9\operatorname {Var}\left[X\right] = 9, find the number nn of observations (the sample size) required in order that with probability less than 5%5\% the mean of the sample differs from the unknown mean μ\mu of XX more,

?
(a)

than 5%5\% of the standard deviation of XX,

(b)

than 5%5\% of μ\mu given that μ>5\mu > 5.

Compare the answers obtained by using Chebyshev's inequality and the CLT.

Problem 264

In a poll designed to estimate the percentage pp of men who support a certain bill, how many men should be questioned in order that, with probability at least 95%95\%, the percentage of the sample differs from pp less,

  1. than 1%1\%,

  2. than 5%5\%, given that,

  3. p<30%p < 30\%, and

  4. pp is completely unknown.

?
Problem 265

Each of the 300 workers of a factory takes his lunch in one of three competing restaurants. How many seats should each restaurant have so that, on average, at most one in 20 customers will remain unseated?

?
Problem 266

The round-off error to the second decimal place has the uniform distribution on the interval (−0.05,0.05)(-0.05, 0.05). What is the probability that the absolute error in the sum of 1,000 numbers is less than 2?

?
Problem 267

The daily income of a card player has uniform distribution in the interval (−40,50)(-40, 50).

?
(a)

What is the probability that he wins more than 500 drachmas in 60 days?

(b)

What amount cc (gain or loss) is such that at most once in 10 times on average the player will have income less than cc during 60 days?

Problem 268

In the game of roulette the probability of winning is 18/3718/37. Suppose that in each game a player earns 1 drachma or loses 1 drachma. How many games must be played daily in a casino in order that the casino earns, with probability 1/21/2, at least 1,000 drachmas daily? Then find the percentage of days on which the casino has a loss.

?
Problem 269

Show that the Γ\Gamma distribution of Exercise 202 tends to the normal distribution as s→∞s \to \infty.

?